OpenStudy (anonymous):

The first term of a sequence is 2 and each successive number is obtained by subtracting 1 from the preceding number and then squaring the result. What is the seventh number in the sequence?

OpenStudy (blockcolder):

Let's write down the terms. :D \[a_1=2\\ a_2=(a_1-1)^2=1^2=1\\ a_3=(a_2-1)^2=0^2=0\\ a_4=(a_3-1)^2=(-1)^2=1\] And a_5 is 0, a_6 is 1, a_7 is....?

OpenStudy (anonymous):

2....??

OpenStudy (blockcolder):

There's a pattern: If you get a 0, the next term is a 1 and if you get a 1, the next term is a 0. Which means that the sequence repeats 0 and 1 ad infinitum. Since a_6 is a 1, a_7 is a ....?

OpenStudy (anonymous):

I don't understand T_T

OpenStudy (anonymous):

2? If it's not two, then I don't know

OpenStudy (blockcolder):

It's a 0. :D

OpenStudy (anonymous):

I HATE SATTTTTTTTTTTTTTT.

OpenStudy (anonymous):

Thank you.

OpenStudy (blockcolder):

Yeah, the questions can be so stupid yet simple. :))